# Monte-Carlo Integration Based Multiple-Scattering Channel Modeling for Ultraviolet Communications in Turbulent Atmosphere

### v1

@@ Abstract
Modeling of multiple-scattering channels in atmospheric turbulence is essential for the performance analysis of long-distance non-line-of-sight (NLOS) ultraviolet (UV) communications. Existing works on the turbulent channel modeling for NLOS UV communications either ignored the turbulence-induced scattering effect or erroneously estimated the turbulent fluctuation effect, resulting in a contradiction with reported experiments. In this paper, we establish a comprehensive multiple-scattering turbulent channel model for NLOS UV communications considering both the turbulence-induced scattering effect and the turbulent fluctuation effect. We first derive the turbulent scattering coefficient and turbulent phase scattering function based on the Booker-Gordon turbulent power spectral density model. Then an improved estimation method is proposed for both the turbulent fluctuation and the turbulent fading coefficient based on the Monte-Carlo integration approach. Numerical results demonstrate that the turbulence-induced scattering effect can always be ignored for typical UV communication scenarios. Besides, the turbulent fluctuation will increase as either the communication distance, the elevation angle, or the divergence angle increases, which is compatible with existing experimental results. Moreover, we find that the probability density of the equivalent turbulent fading for multiple-scattering turbulent channels can be approximated as a Gaussian distribution.
@@ I-A Background and Motivation
The ultraviolet (UV) communication [1, 2] employing the “solar-blind” UV signals ($200$-$280$ nm) as the information carriers has attracted increasing attention in recent decades. Due to its inherent advantages including low background noise, high local security, good adaptivity to extreme weather, and ability of non-line-of-sight (NLOS) links, the UV communication becomes a promising technique for communication scenarios where radio-frequency (RF) communications are impermissible or line-of-sight (LOS) links for traditional free-space optical (FSO) are unavailable [1, 2]. Considerable works have been performed on the NLOS UV channel modeling [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], channel estimation [21, 22, 23, 24, 25, 26], coding and modulation [27, 28, 29, 30], diversity reception [31, 32, 33], duplex and relay techniques [34, 35, 36, 37], and experimental tests [38, 39, 40, 41, 42, 43]. However, most existing works [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] on the channel modeling ignored the impacts of atmospheric turbulence, which cannot be applied to the turbulent cases under either long communication distances or large elevation angles.
The major effects of the atmospheric turbulence on the photon propagating process can be divided into two types: the turbulence-induced scattering effect when a photon interacts with the atmospheric medium [44] and the turbulent fluctuation effect when a photon travels in the atmosphere [45]. In existing works [46, 47, 48, 49, 50, 51, 52], the turbulence-induced scattering effects [46] and the turbulent fluctuation effects [47, 48, 49, 50, 51, 52] were separately studied in the channel modeling of NLOS UV communications. Besides, there exist some contradictions between the existing theoretical turbulent fluctuation model [52] and experimental results [53, 54]. Therefore, a comprehensive turbulent channel model considering both the turbulence-induced scattering effect and the turbulent fluctuation effect is urgently demanded for long-distance NLOS UV communications.
@@ I-B Related Works
The turbulence-induced scattering effect on light wave propagation in the atmosphere was elaborately studied and summarized in [44], where the turbulent atmosphere is modeled as a random continuum with refractive-index varying randomly and continuously in time and space. Some preliminary results on the receiving power and channel impulse response (CIR) of NLOS optical links were derived in [44] based on a single-scattering assumption. The impact of the turbulence-induced scattering effect on the NLOS UV communication was first studied in [46], where the authors claimed that more photons are scattered to the receiver in stronger turbulence for single-scattering cases. However, the impact of the turbulence-induced scattering on the phase scattering function for NLOS UV communications was not clearly studied. Besides, a recent study [55] in underwater optical communication also incorporated the turbulence-induced scattering effect into the scattering-based channel model and demonstrated an addition of more than $5$ dB path loss when the turbulence is considered, which seems opposite to the numerical results obtained in [46]. Therefore, the impact of the turbulence-induced scattering effect on the NLOS UV communication is still not clearly indicated.
@@
The turbulent fluctuation effect on light wave propagating for LOS optical links was elaborately studied and summarized in [45]. The impact of turbulent fluctuation effect on the NLOS UV communications was first studied in a single-scattering model under small common volume assumption [47], where the NLOS link was divided into two LOS links: one from the transmitter to the common volume and another from the common volume to the receiver. When two turbulent fluctuations with log-normal (LN) distribution are introduced into these two LOS links, the authors in [47] demonstrated that the overall turbulent fluctuation can also be approximated as an LN distribution. Based on [47], the bit-error rate performance of NLOS UV communications using this turbulent fluctuation model was studied in [48]. Besides, the turbulent single-scattering model in [47] was extended to narrow beam cases in [49]. A turbulence-induced attenuation effect for LOS links was introduced into the turbulent channel model of NLOS UV communications in [49, 50, 51]. However, the turbulence-induced attenuation considered in LOS laser communications comes from the beam wandering and beam spreading effects [56] caused by the misalignment between the transceivers. Therefore, turbulence-induced attenuation does not apply to NLOS UV links since no alignment is required in NLOS UV communications. Besides, these works [47, 48, 49, 50, 51] are based on the single-scattering model, which cannot be applied to multiple-scattering based long-distance NLOS UV communications.
@@
The turbulent fluctuation model for multiple-scattering was first studied in [52], where the Monte-Carlo based multiple-scattering model [13, 14, 15, 16, 17, 18, 19] is adopted and a Gamma-Gamma (GG) turbulent fading is introduced into each LOS link between two scatters. However, the authors in [52] estimated the turbulent fluctuation by simply calculating the variance of the average receiving power, which cannot really capture the turbulent fluctuation effect. This is because the randomness of the obtained average receiving power in a Monte-Carlo based channel model comes from not only the turbulent fluctuation but also the randomness of the Monte-Carlo process itself [15, 17]. As we will demonstrate in Section IV-C, any attempt to calculate the turbulent fluctuation by estimating the average receiving power of Monte-Carlo based channel models will fail. Based on the same reason, the distribution of the receiving power obtained in [52] cannot characterize the turbulent fading either. Besides, simulation results on the turbulent fluctuation obtained in [52] indicated that the multiple-scattering effect can somehow mitigate the turbulent fluctuation effect, which is opposite to reported experimental results [53, 54]. Therefore, a more sophisticated turbulent fluctuation model is demanded for characterizing the multiple-scattering turbulent channel.
@@ I-C Contributions
Our work aims to establish a multiple-scattering turbulent channel model considering both the turbulence-induced scattering effect and the turbulent fluctuation effect. To achieve this, we first derived the turbulent scattering coefficient and turbulent phase scattering function based on the Booker-Gordon turbulent power spectral density model. Then we proposed an improved estimation method based on a Monte-Carlo integration (MCI) approach for both the turbulent fluctuation and the distribution of the turbulent fading coefficient. The numerical results verified our theoretical analysis and some interesting results were obtained. Specifically, we summarize the main contributions of this work as follows:
  • We established the first comprehensive multiple-scattering turbulent channel model considering both the turbulence-induced scattering effect and fluctuation effect for NLOS UV communications by using an MCI approach.
  • We proposed an improved estimation method for turbulent fluctuation and further proposed an estimation method for the distribution of the turbulent fading coefficient for NLOS UV communications under weak turbulence.
  • We demonstrated that the turbulence-induced scattering effect cannot be ignored only if the turbulent correlation distance approximates the light wavelength under strong turbulence. Besides, for typical UV communication scenarios, we demonstrated that the turbulence-induced scattering effect can always be ignored.
  • We demonstrated that the turbulent fluctuation for NLOS UV communications will increase as either the communication distance, the elevation angle, or the divergence angle increases, which is compatible with existing experimental results.
  • We demonstrated that the distribution of the turbulent fading coefficient for an $n$-order scattering under weak turbulence can be well fitted by a weighted summation of $n$ Gaussian distributions. Besides, the distribution of the overall turbulent fading coefficient for NLOS UV communications can be approximated as a Gaussian distribution.
The rest of this paper is organized as follows. We first introduce the photon propagating model of turbulent channels in Section II. Based on the photon propagating model, we then derive the receiving power of multiple-scattering turbulent channels in Section III. Then we introduce the estimation methods for both the average receiving power, the CIR, the turbulent fluctuation, and the distribution of the turbulent fading coefficient using an MCI approach in Section IV. We then present some numerical results in Section V and conclude our work in Section VI.
@@ II Photon Propagating in Turbulent Channels
The non-turbulent atmosphere can be regarded as a group of particles, which can be modeled as randomly distributed scatters including molecules and aerosols [44]; whereas the turbulent atmosphere is usually regarded as a group of turbulent eddies according to Kolmogorov’s theory of turbulence, which can be modeled as random continuum with refractive-index varying randomly and continuously in time and space [44]. Similar to [55], we consider the scattering effects due to both the random scatterers and the random continuum in this paper. Therefore, the potential scatters include both the particles and the turbulent eddies.
@@ II-A1 Scattering Effect of Particles
When the particle size is much smaller than the light wavelength, the scattering process of a light wave can be modeled as a Rayleigh scattering; whereas when the particle size is comparable to or larger than the light wavelength, the scattering process can be modeled as a Mie scattering [1, 2]. We denote the Rayleigh scattering coefficient and the Mie scattering coefficient by $k_{s}^{ray}$ and $k_{s}^{mie}$, respectively. Then the total scattering coefficient due to the randomly distributed particles is given by $k^{par}_{s}=k_{s}^{ray}+k_{s}^{mie}$.
The scattering strength at different directions around the particle is characterized by the phase scattering function [44]. For UV signals, the phase scattering functions for Rayleigh scattering and Mie scattering are respectively given by [1, 2] $p^{ray}(\theta_{s})=\frac{3[1+3\gamma+(1-\gamma)\cos^{2}\theta_{s}]}{16\pi(1+2\gamma)}$ and $p^{mie}(\theta_{s})=\frac{1-g^{2}}{4\pi}\left[\frac{1}{(1+g^{2}-2g\cos\theta_{s})^{{3}/{2}}}\right.$ $\left.+f\frac{(3\cos^{2}\theta_{s}-1)}{2(1+g^{2})^{{3}/{2}}}\right]$, where $\theta_{s}$ is the scattering (zenith) angle between the incident light direction and the scattering light direction; $\gamma$, $g$, and $f$ are the model parameters. The Rayleigh scattering approximates an isotropic scattering and the Mie scattering is a forward direction dominated scattering [1, 2].
@@ II-A2 Scattering Effect of Turbulence
The turbulence-induced scattering effect for a light propagating in a random continuum can be characterized by the differential cross section per unit volume $\sigma(\bm{i},\bm{o})$ [44], where $\bm{i}$ and $\bm{o}$ are the unit vectors for the incident and the scattering directions, respectively. For a statistically homogeneous and isotropic random continuum and a scalar light wave, the differential cross section $\sigma(\bm{i},\bm{o})$ can be obtained as [44]
$\sigma(\bm{i},\bm{o})=2\pi k^{4}\Phi_{n}(k_{s}),$ (1)
where $\Phi_{n}(k_{s})$ is the power spectral density of the turbulence; $k={2\pi}/{\lambda}$ is the wave number of the incident light; $k_{s}=2k\sin({\theta_{s}}/{2})$ and $\theta_{s}$ is the scattering angle between $\bm{i}$ and $\bm{o}$. Then the scattering coefficient caused by the turbulence can be obtained as [44]
$k_{s}^{tur}=\int_{\Omega}\sigma(\bm{i},\bm{o})\mathrm{d}\omega=4\pi^{2}k^{2}\int_{0}^{2k}\Phi_{n}(k_{s})k_{s}\mathrm{d}k_{s},$ (2)
where $\mathrm{d}\omega$ is the differential solid angle and $\Omega$ is the whole solid angle.
From (1) we can observe that the differential cross section $\sigma(\bm{i},\bm{o})$ is a function of the relative direction between $\bm{i}$ and $\bm{o}$. Then we can rewrite $\sigma(\bm{i},\bm{o})$ as $\sigma(\theta_{s})=2\pi k^{4}\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right)$ and further define the phase scattering function due to the turbulence as
@@
$p^{tur}(\theta_{s})$ $\triangleq\frac{\sigma(\theta_{s})}{\int_{\Omega}\sigma(\theta_{s})\mathrm{d}\omega}=\frac{2\pi k^{4}}{k_{s}^{tur}}\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right).$ (3)
The scattering coefficient $k_{s}^{tur}$ in (2) and the phase scattering function $p^{tur}(\theta_{s})$ in (3) are closely related to the power spectral density $\Phi_{n}(k_{s})$ of the turbulence. However, different assumptions on the turbulence can result in different forms of the power spectral density [44]. Without loss of generality, we adopt the widely used Booker-Gordon model in this paper [44, 46].
The Booker-Gordon model characterizes the turbulent medium by using two quantities, i.e., the variance of the refractive index $\langle n_{1}^{2}\rangle$ and the correlation distance $d_{0}$. Here $n_{1}\triangleq n-\langle n\rangle$ is the fluctuation of the refractive index $n$; $d_{0}$ represents the average size of the turbulent eddies. Then the correlation function of the refractive-index at two positions $\bm{r}_{1}$ and $\bm{r}_{2}$ in Booker-Gordon model is assumed to be [44, 46] $B_{n}(r_{d})\triangleq\langle n_{1}(\bm{r}_{1})n_{1}(\bm{r}_{2})\rangle=\langle n_{1}^{2}\rangle\exp\left({-{r_{d}}/{d_{0}}}\right)$, where $r_{d}=|\bm{r}_{1}-\bm{r}_{2}|$. According to the Wiener-Khinchin theorem [44, 46], the power spectral density $\Phi_{n}(k_{s})$ is the Fourier transform of the correlation function $B_{n}(r_{d})$. For a scalar wave, the power spectral density $\Phi_{n}(k_{s})$ can be obtained as [44, 46]
$\Phi_{n}(k_{s})$ $=\frac{\langle n_{1}^{2}\rangle d_{0}^{3}}{\pi^{2}(1+k_{s}^{2}d_{0}^{2})^{2}}.$ (4)
Then by substituting (4) into (2), we can derive the scattering coefficient of Booker-Gordon model as
$k_{s}^{tur}$ $=\frac{8k^{4}\langle n_{1}^{2}\rangle d_{0}^{3}}{1+4k^{2}d_{0}^{2}}.$ (5)
Besides, when the turbulence is assumed to be isotropic, we have [44, 46] $\langle n_{1}^{2}\rangle=\frac{C_{n}^{2}L_{0}^{{2}/{3}}}{1.91}$, where $L_{0}$ is the outer scale parameter representing the largest turbulent eddy size; $C_{n}^{2}$ is the refractive-index structure parameter representing the strength of the turbulence. The refractive-index structure parameter $C_{n}^{2}$ is defined as the constant coefficient of the refractive-index structure equation [44] $D_{n}(r)\triangleq\langle[n(\bm{r}_{1})-n(\bm{r}_{2})]^{2}\rangle=C_{n}^{2}r^{{2}/{3}}$, where $n(\bm{r}_{i})$ is the refractive-index at position $\bm{r}_{i}$ and $r=|\bm{r}_{1}-\bm{r}_{2}|$ is the distance between two positions $\bm{r}_{1}$ and $\bm{r}_{2}$. The refractive-index structure equation also provides a way for measuring $C_{n}^{2}$ in practical implementations. Typical values of $L_{0}$ can vary from the magnitude of $10^{0}$ m to $10^{3}$ m; typical values of $C_{n}^{2}$ are at the magnitude of $10^{-15}$ $\text{m}^{-2/3}$ and $C_{n}^{2}$ can vary from $10^{-17}$ for weak turbulence to $10^{-13}$ for strong turbulence [44, 46]. Then by substituting $\langle n_{1}^{2}\rangle$ into (5), we can further obtain
$k_{s}^{tur}=\frac{8k^{4}C_{n}^{2}L_{0}^{{2}/{3}}d_{0}^{3}}{1.91(1+4k^{2}d_{0}^{2})}.$ (6)
From the expression of $k_{s}^{tur}$ in (6), we can observe that $k_{s}^{tur}\propto C_{n}^{2}$. Besides, when $k^{2}d_{0}^{2}\ll 1$, we have $k_{s}^{tur}\approx\frac{8k^{4}C_{n}^{2}L_{0}^{{2}/{3}}d_{0}^{3}}{1.91}$, which indicates that $k_{s}^{tur}\propto d_{0}^{3}$ when the correlation distance $d_{0}$ is much smaller than the wavelength $\lambda$. In this case, we also have $k_{s}^{tur}\propto 1/\lambda^{4}$, which coincides with the Rayleigh scattering. When $k^{2}d_{0}^{2}\gg 1$, we have $k_{s}^{tur}\approx\frac{2k^{2}C_{n}^{2}L_{0}^{{2}/{3}}d_{0}}{1.91}$, which indicates that $k_{s}^{tur}\propto d_{0}$ when the correlation distance $d_{0}$ is much larger than the wavelength $\lambda$.
@@
Substituting (4) and (5) into (3), we can obtain the phase scattering function for the Booker-Gordon model as
@@
From the phase scattering function $p^{tur}(\theta_{s})$ in (7), we can observe that the phase scattering function depends only on the wave number $k$ and the correlation distance $d_{0}$. Besides, when $k^{2}d_{0}^{2}\ll 1$, we have $p^{tur}(\theta_{s})\approx 1/4\pi$, which indicates that the scattering will approach an isotropic scattering when the correlation distance $d_{0}$ is much smaller than the wavelength $\lambda$. This isotropic scattering property is similar to the Rayleigh scattering. When $k^{2}d_{0}^{2}\gg 1$, the majority of the scattering intensity happens at small scattering angles. This indicates that the scattering is forward direction dominated when the correlation distance $d_{0}$ is much larger than the wavelength $\lambda$, which is similar to the Mie scattering.
Combining the coefficients $k_{s}^{ray}$, $k_{s}^{mie}$, and $k_{s}^{tur}$, we can obtain the total scattering coefficients as
$k_{s}^{tot}=k_{s}^{ray}+k_{s}^{mie}+k_{s}^{tur}.$ (8)
Similarly, combining the phase scattering functions $p^{ray}(\theta_{s})$, $p^{mie}(\theta_{s})$, and $p^{tur}(\theta_{s})$, we can obtain the total phase scattering function as a weighted function of $p^{ray}(\theta_{s})$, $p^{mie}(\theta_{s})$, and $p^{tur}(\theta_{s})$, i.e.,
$p^{tot}(\theta_{s})=\frac{k_{s}^{ray}}{k_{s}^{tot}}p^{ray}(\theta_{s})+\frac{k_{s}^{mie}}{k_{s}^{tot}}p^{mie}(\theta_{s})+\frac{k_{s}^{tur}}{k_{s}^{tot}}p^{tur}(\theta_{s}).$ (9)

### v3

# Turbulent Multiple-Scattering Channel Modeling for Ultraviolet Communications: A Monte-Carlo Integration Approach

@@ Abstract
Modeling of multiple-scattering channels in atmospheric turbulence is essential for the performance analysis of long-distance non-line-of-sight (NLOS) ultraviolet (UV) communications. Existing works on the turbulent channel modeling for NLOS UV communications either focused on single-scattering cases or estimate the turbulent fluctuation effect in an unreliable way based on Monte-Carlo simulation (MCS) approach. In this paper, we establish a comprehensive turbulent multiple-scattering channel model by using a more efficient Monte-Carlo integration (MCI) approach for NLOS UV communications, where both the scattering, absorption, and turbulence effects are considered. Compared with the MCS approach, the MCI approach is more interpretable for estimating the turbulent fluctuation. To achieve this, we first introduce the scattering, absorption, and turbulence effects for NLOS UV communications in turbulent channels. Then we propose the estimation methods based on MCI approach for estimating both the turbulent fluctuation and the distribution of turbulent fading coefficient. Numerical results demonstrate that the turbulence-induced scattering effect can always be ignored for typical UV communication scenarios. Besides, the turbulent fluctuation will increase as either the communication distance increases or the zenith angle decreases, which is compatible with existing experimental results and also with our experimental results. Moreover, we demonstrate numerically that the distribution of the turbulent fading coefficient for UV multiple-scattering channels under all turbulent conditions can be approximated as log-normal distribution; and we also demonstrate both numerically and experimentally that the turbulent fading can be approximated as a Gaussian distribution under weak turbulence.
@@ I-A Background and Motivation
The ultraviolet (UV) communication [yuan2016review] employing the “solar-blind” UV signals (200200-$280$ nm) as the information carriers has attracted increasing attention in recent decades. Due to its inherent advantages including low background noise, high local security, good adaptivity to extreme weather, and ability of non-line-of-sight (NLOS) links, the UV communication becomes a promising technique for communication scenarios where radio-frequency (RF) communications are impermissible or line-of-sight (LOS) links for traditional free-space optical (FSO) are unavailable [yuan2016review]. In recent decade, considerable works have been performed on the NLOS UV channel modeling [yuan2016integral, wu2019single, shen2019modeling, cao2021single, cao2022single, ding2009modeling, drost2011uv, ding2010path, yuan2019importance, yuan2019monte, shan2020modeling], channel estimation [gong2015channel, gong2016optical, wei2018simultaneous, hu2020non, shen2021lmmse], coding and modulation [cao2023performance, cao2023power], diversity reception [qin2017noncoplanar, yuan2023simo, wang2023mimo], duplex and relay techniques [ardakani2017performance, gong2018full, wang2023nlos], and experimental tests [sun201771, wang20181mbps, alkhazragi2020gbit]. However, most existing works on the channel modeling ignored the impacts of atmospheric turbulence, which cannot be applied to the turbulent cases under either long communication distances or large elevation angles.
In existing literatures, the major effects of the atmospheric turbulence on the photon propagating process can be divided into two types: the turbulence-induced scattering effect when a photon interacts with the atmospheric medium [ishimaru2005theory] and the turbulent fluctuation effect when a photon travels in the atmosphere [andrews2005laser]. In existing works, the turbulence-induced scattering effects and the turbulent fluctuation effects were separately studied in the channel modeling of NLOS UV communications. Besides, existing works on turbulent effects in NLOS UV communications either focused on single-scattering cases or estimate the turbulent fluctuation based on a Monte-Carlo simulation (MCS) approach. As it is demonstrated in [yuan2019importance, yuan2019monte], the Monte-Carlo integration (MCI) approach can provide a more efficient modeling for multiple scattering channels. Besides, as we will demonstrate later in this paper, the MCI approach can also provide a more interpretable logic for estimating the turbulent fluctuation. Therefore, a comprehensive turbulent multiple scattering channel modeling based on MCI approach considering both the scattering, absorption, and turbulence effects is urgently demanded for long-distance NLOS UV communications.
@@ I-B Related Works
The turbulence-induced scattering effect on light wave propagation in the atmosphere was elaborately studied and summarized in [ishimaru2005theory], where the turbulent atmosphere is modeled as a random continuum with refractive-index varying randomly and continuously in time and space. Some preliminary results on the receiving power and channel impulse response (CIR) of NLOS optical links were derived in [ishimaru2005theory] based on a single-scattering assumption. The impact of the turbulence-induced scattering effect on the NLOS UV communication was first studied in [xiao2013non]. However, the impact of the turbulence-induced scattering on the phase scattering function for NLOS UV communications was not clearly studied. Besides, a recent study [xu2022improvement] in underwater optical communication also incorporated the turbulence-induced scattering effect into the scattering-based channel model and demonstrated an addition of more than $5$ dB path loss when the turbulence is considered, which seems opposite to the numerical results obtained in [xiao2013non]. Therefore, the impact of the turbulence-induced scattering effect on the NLOS UV communication is still not clearly indicated.
@@
The turbulent fluctuation effect on light wave propagating for LOS optical links was elaborately studied and summarized in [andrews2005laser]. The impact of turbulent fluctuation effect on the NLOS UV communications was first studied in a single-scattering model under small common volume assumption [ding2011turbulence], where the NLOS link was divided into two LOS links: one from the transmitter to the common volume and another from the common volume to the receiver. When two turbulent fluctuations with log-normal (LN) distribution are introduced into these two LOS links, the authors in [ding2011turbulence] demonstrated that the overall turbulent fluctuation can also be approximated as an LN distribution. Based on [ding2011turbulence], the bit-error rate performance of NLOS UV communications using this turbulent fluctuation model was studied in [liu2015performance]. Besides, the turbulent single-scattering model in [ding2011turbulence] was extended to narrow beam cases in [shan2020single]. A turbulence-induced attenuation effect for LOS links was introduced into the turbulent channel model of NLOS UV communications in [shan2020single, zuo2012effect, xiao2012non]. However, the turbulence-induced attenuation considered in LOS laser communications comes from the beam wandering and beam spreading effects [kaushal2017free] caused by the misalignment between the transceivers. Therefore, turbulence-induced attenuation cannot directly be applied to NLOS UV links since no alignment is required in NLOS UV communications. Recently, a turbulent channel modeling based on equivalent scattering points approach was proposed in [chu2024turbulent], where a Gaussian distribution on the turbulent fading coefficient was obtained. However, these works [ding2011turbulence, liu2015performance, shan2020single, zuo2012effect, xiao2012non, chu2024turbulent] are based on the single-scattering model, which cannot be applied to multiple-scattering based long-distance NLOS UV communications.
@@
The turbulent fluctuation model for multiple-scattering was first studied in [wang2013characteristics] based on an MCS approach [drost2011uv], where a Gamma-Gamma (GG) turbulent fading is introduced into each LOS link between two scatters. However, the authors in [wang2013characteristics] estimated the turbulent fluctuation by simply calculating the variance of the average receiving power. As we will demonstrate in Section IV-B, the attempt to calculate the turbulent fluctuation by estimating the average receiving power of Monte-Carlo based channel models is unreliable due to the properties of Monte-Carlo process. Besides, it was demonstrated that the MCI approaches enjoy more efficient modeling compared with the MCS approaches. Therefore, a more sophisticated turbulent fluctuation model based on the MCI approach is demanded for characterizing the turbulent multiple-scattering channel.
@@ I-C Contributions
Our work aims to establish a turbulent multiple-scattering channel model based on an MCI approach for NLOS UV communications, considering both scattering, absorption, and turbulence effects. To achieve this, we first introduce the scattering, absorption, and turbulence effects of NLOS UV communications in turbulent channels. Then we proposed an improved estimation method based on an MCI approach for both the turbulent fluctuation and the distribution of the turbulent fading coefficient. Both numerical simulation and experimental measurements are conducted to verified our theoretical analysis. Specifically, we summarize the main contributions of this work as follows:
  • We established the first turbulent multiple-scattering channel model based on an MCI approach for NLOS UV communications, where the scattering, absorption, and turbulence effects are considered.
  • We proposed an improved estimation method for turbulent fluctuation and further proposed an estimation method for the distribution of the turbulent fading coefficient for NLOS UV communications.
  • We demonstrated that the turbulence-induced scattering effect can always be ignored for typical UV communication scenarios; and we demonstrated that the turbulent fluctuation for NLOS UV communications will increase as either the communication distance increases or the zenith angle decreases, which is compatible with existing experimental results and also with our experimental results.
  • We demonstrated numerically that the distribution of the turbulent fading coefficient for NLOS UV communications in all turbulent conditions can be approximated by LN distributions; and we also demonstrated both numerically and experimentally that the turbulent fading under weak turbulence can be approximated as a Gaussian distribution.
The rest of this paper is organized as follows. We first introduce the photon propagating model of turbulent channels in Section II. Based on the photon propagating model, we then derive the receiving power of multiple-scattering turbulent channels in Section III. Then we introduce the estimation methods based on an MCI approach for both the average receiving power, the turbulent fluctuation, and the distribution of the turbulent fading coefficient in Section IV. We then present some numerical and experimental results in Section V and conclude our work in Section VI.
@@ II Photon Propagating in Turbulent Channels
The non-turbulent atmosphere can be regarded as a group of particles, which can be modeled as randomly distributed scatters including molecules and aerosols [ishimaru2005theory]; whereas the turbulent atmosphere is usually regarded as a group of turbulent eddies according to Kolmogorov’s theory of turbulence, which can be modeled as random continuum with refractive-index varying randomly and continuously in time and space [ishimaru2005theory]. Similar to [xu2022improvement], we consider the scattering effects due to both the random scatterers and the random continuum in this paper. Therefore, the potential scatters include both the particles and the turbulent eddies.
@@ II-A1 Scattering Effect of Particles
When the particle size is much smaller than the light wavelength, the scattering process of a light wave can be modeled as a Rayleigh scattering; whereas when the particle size is comparable to or larger than the light wavelength, the scattering process can be modeled as a Mie scattering [yuan2016review]. We denote the Rayleigh scattering coefficient and the Mie scattering coefficient by $k_{s}^{ray}$ and $k_{s}^{mie}$, respectively. Then the total scattering coefficient due to the randomly distributed particles is given by
$k^{par}_{s}=k_{s}^{ray}+k_{s}^{mie}.$ (1)
The scattering strength at different directions around the particle is characterized by the phase scattering function [ishimaru2005theory]. For UV signals, the phase scattering functions for Rayleigh scattering and Mie scattering are given by [yuan2016review]
$\begin{cases}p^{ray}(\theta_{s})\!=\!\frac{3[1+3\gamma+(1-\gamma)\cos^{2}\theta_{s}]}{16\pi(1+2\gamma)},\\ p^{mie}(\theta_{s})\!=\!\frac{1-g^{2}}{4\pi}\!\!\left[\frac{1}{(1+g^{2}\!-\!2g\cos\theta_{s})^{{3}/{2}}}\!+\!\frac{f(3\cos^{2}\theta_{s}\!-\!1)}{2(1+g^{2})^{{3}/{2}}}\right],\end{cases}$ (2)
where $\theta_{s}$ is the scattering (zenith) angle between the incident light direction and the scattering light direction; $\gamma$, $g$, and $f$ are the model parameters. The Rayleigh scattering approximates an isotropic scattering and the Mie scattering is a forward direction dominated scattering [yuan2016review].
@@ II-A2 Scattering Effect of Turbulence
The turbulence-induced scattering effect for a light propagating in a random continuum can be characterized by the differential cross section per unit volume $\sigma(\bm{i},\bm{o})$ [ishimaru2005theory], where $\bm{i}$ and $\bm{o}$ are the unit vectors for the incident and the scattering directions, respectively. For a statistically homogeneous and isotropic random continuum and a scalar light wave, the differential cross section $\sigma(\bm{i},\bm{o})$ can be obtained as $\sigma(\bm{i},\bm{o})=2\pi k^{4}\Phi_{n}(k_{s})$, where $\Phi_{n}(k_{s})$ is the power spectral density of the turbulence; $k={2\pi}/{\lambda}$ is the wave number of the incident light; $k_{s}=2k\sin({\theta_{s}}/{2})$ and $\theta_{s}$ is the scattering angle between $\bm{i}$ and $\bm{o}$. We can observe that the differential cross section $\sigma(\bm{i},\bm{o})$ is a function of the relative direction between $\bm{i}$ and $\bm{o}$. Then we can rewrite $\sigma(\bm{i},\bm{o})$ as $\sigma(\theta_{s})=2\pi k^{4}\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right)$ and further define the phase scattering function due to the turbulence as [ishimaru2005theory]
@@
$p^{tur}(\theta_{s})$ $\triangleq\frac{\sigma(\theta_{s})}{\int_{\Omega}\sigma(\theta_{s})\mathrm{d}\omega}=\frac{\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right)}{\int_{\theta_{s}}\Phi_{n}\left(2k\sin({\theta_{s}}/{2})\right)\mathrm{d}\theta_{s}}.$ (3)
Without loss of generality, we adopt the widely used Booker-Gordon model [ishimaru2005theory, xiao2013non] with $\Phi_{n}(k_{s})=\frac{\langle n_{1}^{2}\rangle d_{0}^{3}}{\pi^{2}(1+k_{s}^{2}d_{0}^{2})^{2}}$, where $\langle n_{1}^{2}\rangle$ is the variance of refractive index; $d_{0}$ is the average size of turbulent eddies. When the turbulence is assumed to be isotropic, we have [ishimaru2005theory, xiao2013non] $\langle n_{1}^{2}\rangle=\frac{C_{n}^{2}L_{0}^{{2}/{3}}}{1.91}$, where $L_{0}$ is the outer scale parameter representing the largest turbulent eddy size; $C_{n}^{2}$ is the refractive-index structure parameter representing the strength of the turbulence.
@@
Then substituting $\Phi_{n}(k_{s})$ into (3), we can obtain the phase scattering function due to turbulence as
@@
Because for optical wavebands including UV wavebands, we always have $\lambda\ll d_{0}$, i.e., $k^{2}d_{0}^{2}\gg 1$. This indicates that the scattering happens always at forward direction, which is equivalent to a rectilinear propagation effect. Therefore, the turbulence induced scattering effect can be omitted since it has no true effect on the propagating direction of the light.
At last, we can obtain the scattering coefficient in turbulent channels as
$k_{s}^{tot}=k_{s}^{par}=k_{s}^{ray}+k_{s}^{mie}.$ (5)
Similarly, combining the phase scattering functions $p^{ray}(\theta_{s})$ and $p^{mie}(\theta_{s})$, we can obtain the total phase scattering function in turbulent channel as
$p^{tot}(\theta_{s})=\frac{k_{s}^{ray}}{k_{s}^{tot}}p^{ray}(\theta_{s})+\frac{k_{s}^{mie}}{k_{s}^{tot}}p^{mie}(\theta_{s}).$ (6)